OpenStudy (anonymous):

Solve the inequality. Write answer in interval notation. x^2+4x-2>0

OpenStudy (anonymous):

\[x=[-4\pm \sqrt{4^2-4(1)(-2)}]/2=-2\pm \sqrt{16+8}=-2\pm 2\sqrt{6}\] That would satisfy x^2+4x-2=0 so plug in values greater or less than your found x's and see which makes the inequality true. For example plug in number less than -2+2sqrt(6) we can use -3+2sqrt(6) plug it into x^2+4x-2 and you get 9.202 which is greater than 0, thus your solution includes x<-2+2sqrt(6) Do you follow?